OpenStudy (anonymous):

Arc Lengths: Find the exact length of the polar curve r= 5sin(theta), 0<= theta <= pi/3.

OpenStudy (anonymous):

LOLOLLOLOL

OpenStudy (anonymous):

do you know the arclength formula?

OpenStudy (anonymous):

LOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOLLOLOL

OpenStudy (anonymous):

bro are you okay?

OpenStudy (anonymous):

\[L=\int\limits_{0}^{pi/3} \sqrt{1+[dy/dx]^2} d\]

OpenStudy (anonymous):

No, he's not. Im gonna report him cause he's annoying me.

OpenStudy (anonymous):

yeah you got it in this case instead of dy/dx we will have dr/dtheta okay

OpenStudy (anonymous):

and lol do it up so anyways

OpenStudy (anonymous):

derivative of r with respect to theta is 5*cos(theta) alright now let us put this into our formula KEEP in mind for these problems go inside out like do the derivative then square it then add 1 then square root it then take the integral alright ok so anyways now we got 1+ 25cos^2theta

OpenStudy (anonymous):

now we have to take the square root of that and now we apply trignometric substitution here

OpenStudy (anonymous):

to find the main integral

OpenStudy (anonymous):

trignometric substitution !?!? Oh dear. I especially have issues with those.

OpenStudy (anonymous):

yeah because there is no other way to actually integrate this :) so make the right substitution and work from there or do you want help with that?

OpenStudy (anonymous):

Ummmmmmmmm, im gonna need help. You're help in two ways since I also need help with trig sub. I know that 1+cos^2 = sin^2. and something something...

OpenStudy (anonymous):

no 1-cos^2x = sin^2x

OpenStudy (anonymous):

heck, smh

OpenStudy (anonymous):

to solve this integral here take a look at this page for me it is really usefull it helped me all the way till 4th year uni http://www.stewartcalculus.com/data/CALCULUS%20Concepts%20and%20Contexts/upfiles/3c3-TrigonometSubstitu_Stu.pdf

OpenStudy (anonymous):

go down a little to see trignometric substitution table

OpenStudy (anonymous):

i should use...the tan one?

OpenStudy (anonymous):

exactly :) you got it where your x=5cosx alright :)

OpenStudy (anonymous):

Ok, I tried and failed horribly, i am so sorry.

OpenStudy (anonymous):

\[\sqrt{5^2-(5\cos(\theta))^2} = \sqrt{25-25\cos^2(\theta)} = \sqrt{5(1-\cos^2(\theta)} =\sqrt{5^2}\sqrt{1-\cos^2(\theta)} \]

OpenStudy (anonymous):

\[= 5 \sqrt{\sin^2(\theta)} = 5 \sin(\theta)\]

OpenStudy (anonymous):

no that is wrong my friend

OpenStudy (anonymous):

good try though you were suppose to use tan remember from that table that table basically tells you if you have something of this format use this and apply this identity right?

OpenStudy (anonymous):

yeaa. But because I have cos and sin, i got confused i guess

OpenStudy (anonymous):

o wow which year are you in by any chance?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+sqrt%281+%2B+25cosx%5E2%29 according to wolfram you cannot integrate even with trig substituion you cannot apply it actually ah mann

OpenStudy (anonymous):

This is my second year of college, calc 2

OpenStudy (anonymous):

How'd u make your wolfram link so short?

OpenStudy (anonymous):

what do u meann

OpenStudy (anonymous):

nvm. When i used wolfram, they said the answe was 4.4... integral_0^(pi/3) sqrt(1+25 cos^2(t)) dt = sqrt(26) E(pi/3|25/26)~~4.45953

OpenStudy (anonymous):

This is actually a test question. I dont see how he expected us to know how to do this

OpenStudy (anonymous):

I realized the issue, lol. this is a polar curve., so \[L= \int\limits_{0}^{pi/3} \sqrt{(r)^2 + [dr/dt]^2} dt\]